OpenStudy (anonymous):

How do you maximize Q=x^2y, where x and y are positive and x+y=60?

OpenStudy (anonymous):

start by noting that if \(x+y=60\) then \(y=60-x\) and maximize \(x^2(60-x)\) on \((0,\infty)\)

OpenStudy (anonymous):

go to \(f(x)=60x^2-x^3\) take the derivative, set it equal to zero to find the critical points etc

OpenStudy (anonymous):

yeah i made a mistake...

OpenStudy (anonymous):

maximize on \((0,60)\)

terenzreignz (terenzreignz):

For a nice few minutes there, I read it as \[\huge x^{2y}\] >.<

OpenStudy (anonymous):

have fun with that one!!